# Counting-Stars (★): A Multi-evidence, Position-aware, and Scalable Benchmark for Evaluating Long-Context Large Language Models

### v3

@@ Abstract
While recent research endeavors have focused on developing Large Language Models (LLMs) with robust long-context capabilities, due to the lack of long-context benchmarks, relatively little is known about how well the performance of long-context LLMs. To address this gap, we propose a multi-evidence, position-aware, and scalable benchmark for evaluating long-context LLMs, named Counting-Stars, which evaluates long-context LLMs by using two tasks: multi-evidence acquisition and multi-evidence reasoning. Based on the Counting-Stars test, we conduct experiments to evaluate long-context LLMs (i.e., GPT-4 Turbo, Gemini 1.5 Pro, Claude3 Opus, GLM-4, and Moonshot-v1). Experimental results demonstrate that Gemini 1.5 Pro achieves the best overall results, while the performance of GPT-4 Turbo is the most stable across various tasks. Furthermore, our analysis of these LLMs, which are extended to handle long-context scenarios, indicates that there is potential for improvement as the length of the input context and the intricacy of the tasks are increasing. The code and data are released  [^1] https://github.com/nick7nlp/Counting-Stars.
@@ A Multi-evidence, Position-aware, and Scalable Benchmark for
Mingyang Song, Mao Zheng, Xuan Luo Tencent MLPD nickmysong@tencent.com
@@ 1 Introduction
A popular benchmark for whether LLMs are able to utilize long context is the needle-in-a-haystack [^2] https://github.com/gkamradt/LLMTest_NeedleInAHaystack, requiring LLMs to precisely find and recite the evidence in a sentence where the sentence is inserted in an arbitrary location of the long context, which evaluates the capability of evidence acquisition of the long-context LLMs. However, many recently released long-context LLMs adopt the needle-in-a-haystack to evaluate the ability of long-context handling and achieve nearly perfect performance, making the needle-in-a-haystack impossible to distinguish the gaps between the long-context LLMs. This not only demonstrates the improvement of the recent long-context LLMs but also denotes that the needle-in-a-haystack is too simple to further test the capability of the long-context LLMs.
To mitigate the shortcomings of existing benchmarks, in this paper, we propose a multi-evidence, position-aware, and scalable benchmark for evaluating long-context LLMs as a novel benchmark, named Counting-Stars. As the name suggests, the Counting-Stars test refers to asking LLMs to collect the numbers of stars from multiple sentences describing the number of stars counted by the little penguin inserted in the long context and then summarize into a specified answer. Through the Counting-Stars, we expect to evaluate the long context capabilities of multi-evidence acquisition and multi-evidence reasoning of LLMs. More specifically, the former focuses on testing the capability of LLMs to retrieve evidence at different positions within the long context, which can more clearly reflect the quality of long-context modeling. When collecting evidence, reasoning is often required to ensure that the evidence gathered supports the correct answer to the question. Therefore, the latter evaluates the LLM’s ability to filter out noise or incorrect information when retrieving information and the model’s reasoning ability at different positions within the long context. Generally speaking, the latter is definitely more challenging than the former. In other words, the former can be treated as making LLMs distinguish between long contexts and inserted sentences (similar to the needle-in-a-haystack), while the latter involves distinguishing evidence within each inserted sentence.
@@
Multi-evidence Acquisition
@@
Table 1: Two task descriptions of the Counting-Stars test.
@@ 2 Counting-Stars (★)
Multi-evidence. In long-context scenarios, answering a question may require collecting lots of evidence from different positions of the long context. Hence, it is necessary to verify the capability of LLMs to collect lots of evidence from a long context at once. In addition, to the best of our knowledge, the Counting-Stars is the first long-context benchmark to increase significantly the number of pieces of evidence (i.e., increase to $32$, $64$, $128$, $256$, $512$, and even $1024$).
@@
In general, the capacity for a long-context LLM to do human textual instructions largely depends on its multi-evidence acquisition ability. Moreover, an indispensable ability of LLMs extends beyond mere essential evidence collection to encompass reasoning based on the collected evidence. Therefore, the Counting-Stars mainly evaluates the long-context capability of LLMs from two perspectives, i.e., long-context multi-evidence acquisition and long-context multi-evidence reasoning. Expressly, this test can be understood as asking LLMs to find and remember all the sentences in the long text that describe the little penguin counting stars and organize them into a list to return the final answer. All sentences are prior inserted into a long text at the same interval. In addition, the used long text can be any text data that is not related to the sentences describing the little penguin counting stars, such as The Story of the Stone [^3] The Story of the Stone, is an 18th-century Chinese novel authored by Cao Xueqin, considered to be one of the Four Great Classical Novels of Chinese literature. for the Chinese version of the Counting-Stars and Paul Graham Essays [^4] The English context data used in this paper is similar to the needle-in-a-haystack. for the English version of the Counting-Stars. Next, we will introduce the Counting-Stars test in detail.
@@
2.1 Long-Context Multi-evidence Acquisition
Multi-evidence acquisition refers to the capability of distinguishing and collecting critical information framed within intricate and long textual data, which bottlenecks the performance of LLMs in synthesizing contextualized knowledge to execute various tasks, from answering multi-document questions to executing complex human instructions. Furthermore, maintaining a comprehensive and accurate grasp of the input text becomes increasingly challenging as the context length increases. Therefore, in the Counting-Stars test, the first task is to examine the multi-evidence acquisition ability of long-context LLMs, as illustrated in Table 1 (named as Long-Context Multi-evidence Acquisition). In multi-evidence acquisition, all sentences describing the little penguin counting stars are designed as "The little penguin counted {number1}". Here, {number1} indicates the number of stars the little penguin counted. Concretely, we randomly generated all the numbers of stars as {number1, number2, …} because we found that LLMs easily slack off if a sequence of numbers is continuous or regular. In this task, we hope that LLMs collect all the numbers of stars the little penguin counted and summarize them in a list.
@@ 2.3 Scalable Test Setting
Multi-evidence Acquisition (ZH) 0.697 0.663 0.775 0.807 0.788 0.698 0.682 0.606
Multi-evidence Acquisition (EN) 0.718 0.662 0.833 0.705 - - 0.389 0.559
@@
Table 3: The overall performance of LLMs on the Counting-Stars-(32) test.
Models Multi-evidence Acquisition (ZH)
@@
Table 4: The performance of LLMs on the Chinese version of the Counting-Stars-(32)-(Multi-evidence Acquisition).
Models Multi-evidence Acquisition (EN) Multi-evidence Reasoning (ZH) Multi-evidence Reasoning (EN)
@@
Table 5: The performance of LLMs on the English version of the Counting-Stars-(32)-(Multi-evidence Acquisition) as well as the Chinese and English versions of the Counting-Stars-(32)-(Multi-evidence Reasoning).
@@ 3.1 Baselines and Experimental Settings
Before evaluating, we truncate the results to the length of the ground truth and eliminate duplicates. For a piece of test data, the prediction results are evaluated starting from the first number of stars, that is, {number1}, {number2}, …, {numberM}. For the first task, if the results contains {number1}, it gets a score of $1$; if it doesn’t, it gets 00. For the second task, when the results contains only {number1}, the score is $1$; if it also contains {wrong number1}, the score is $0.5$; if the value only contains {wrong number1}, the score is $0.25$, and if both If not found, the score is 00. Generally, text evaluation is usually more complex, so the evidence to be collected in the Counting-Stars is all numerical, making it more straightforward to evaluate.
Figure 2: Visualization of the results on the Chinese version of the Counting-Stars-32-(Multi-evidence Acquisition).
@@ 3.2 Overall Performance
Table 3 present the testing performance of GPT-4 Turbo, Claude3 Opus, Gemini 1.5 Pro, GLM-4, Moonshot-v1 on the Chinese and English versions of the Counting-Stars-(32)-(Multi-evidence Acquisition) and Counting-Stars-(32)-(Multi-evidence Reasoning) tests. Overall, Claude3 Opus achieves the best performance on the Chinese version of the Counting-Stars-(32)-(Multi-evidence Acquisition), Gemini 1.5 Pro obtains the best performance on the English version of the Counting-Stars-(32)-(Multi-evidence Acquisition) and the Chinese version of the Counting-Stars-(32)-(Multi-evidence Reasoning), and GPT-4 Turbo obtains the best performance on the English version of the Counting-Stars-(32)-(Multi-evidence Reasoning). Although these long-context LLMs have achieved nearly perfect performance on the needle-in-a-haystack task, they still perform poorly on the Counting-Stars test, which indicates that the needle-in-a-haystack is too simple to truly show the capabilities of LLMs in processing long texts.
The multi-evidence reasoning task necessitates that LLMs engage in acquiring and reasoning multiple pieces of evidence simultaneously, which is more complex than the multi-evidence acquisition task. This task requires LLMs to sift through and exclude inaccurate evidence while gathering information from a long context to answer questions. As indicated by the data in Table 3, each LLM performs not well enough. Notably, in contrast to GPT-4 Turbo and Claude3 Opus, Gemini 1.5 Pro stands out for having gathered virtually no incorrect information, as shown in Figure 3.
From Table 4 and Table 5, it can be observed that all LLMs are capable of achieving higher scores in short-context scenarios, which confirms that the Counting-Stars is reasonable and can be accomplished by LLMs. However, as the context length increases, the performance of all models shows a downward trend. Among them, GPT-4 Turbo’s performance is relatively stable. In addition, GLM-4 has obtained surprising results under the 32K context length of the Chinese version of the Counting-Stars-(32)-(Multi-evidence Acquisition).
@@ 4.1 The Length-Stability Dilemma
However, based on the experiments in this paper, we cannot yet determine the specific reasons, which is also a goal that the next version of Counting-Stars aims to achieve. We consider that the most intuitive idea behind this is that the long-context capability of LLMs is still relatively weak, so in the case of limited resources, part of stability needs to be sacrificed. Addressing this issue may aid researchers in better analyzing and enhancing the long-context modeling capabilities of LLMs, benefiting specific NLP tasks, such as multi-document question answering. By the way, stability refers to the understanding and reasoning ability of LLMs when handling different long contexts, which is more crucial than the length of context processing.
@@ 5 Related Work
The Counting-Stars evaluates the capabilities of multi-evidence acquisition and reasoning of LLMs, which should be more noteworthy in the long context modeling of LLMs, as reflected in tasks such as multi-document question answering and summarization. Concretely, the former primarily evaluates the capability of LLMs to collect multiple pieces of evidence simultaneously (distinguishing between long context and inserted sentences), while the latter tests the ability of LLMs to gather and reason various pieces of evidence at the same time correctly, that is, reasoning is required when collecting information (distinguishing between correct and incorrect evidence in inserted sentences). To the best of our knowledge, the Counting-Stars is the first scalable long-context benchmark to ask LLMs to simultaneously differentiate between correct and incorrect evidence in each inserted sentence.
@@ 6 Conclusion
We introduce a multi-evidence, position-aware, and scalable benchmark for evaluating long-context LLMs. Based on the Counting-Stars, we conduct intriguing analyses on the behavior of LLMs, including the length-stability dilemma and the absence of the lost-in-the-middle phenomenon. Our testing and analysis provide valuable insights into LLMs when dealing with long contexts, which can inform and guide future research endeavors.
@@ 7 Limitations
In the future, we will evaluate more famous LLMs on the Counting-Stars, such as Llama3888https://github.com/meta-llama/llama3, Mistral [^9] https://mistral.ai/news/la-plateforme/, Llama2 Touvron et al. (2023), Yi AI et al. (2024), Mixtral Jiang et al. (2024), Command-R [^10] https://docs.cohere.com/docs/command-r#model-details, LongLoRA Chen et al. (2024), LongAlpacaChen et al. (2024), LWM Liu et al. (2024), Jamba Lieber et al. (2024), QWen Bai et al. (2023a), DeepSeek-V2 DeepSeek-AI (2024), etc.
@@
Initially, the Counting-Stars is designed to require LLMs to count the total number of stars in all sentences inserted in the long context, which aims to test the multi-evidence acquisition of LLMs from a long dependency perspective. However, we find that if LLMs are required to calculate the total number of stars, they usually perform poorly. Specifically, we analyze the reasons for the bad performance, which mainly include three points:

### v4

@@ Abstract
Despite recent efforts to develop large language models with robust long-context capabilities, the lack of long-context benchmarks means that relatively little is known about their performance. To alleviate this gap, in this paper, we propose Counting-Stars, a multi-evidence, position-aware, and scalable benchmark designed to evaluate the multi-evidence retrieval capabilities of long-context LLMs. Counting-Stars comprises two counting-based multiple pieces of evidence retrieval tasks: searching and reasoning. Using Counting-Stars, we conducted experiments to evaluate several long-context LLMs, including GPT-4 Turbo, Gemini 1.5 Pro, Claude3 Opus, GLM-4, and Moonshot-v1. Extensive experimental results demonstrate that Gemini 1.5 Pro achieves the best overall results, while GPT-4 Turbo exhibits the most stable performance across various tasks. Furthermore, our analysis of these LLMs, which have been extended to handle long-context scenarios, indicates that significant room for improvement remains as the length of the input context and the complexity of the tasks increase. The code and data are publicly accessible here [^1] https://github.com/nick7nlp/Counting-Stars.
@@ A Multi-evidence, Position-aware, and Scalable Benchmark for
Mingyang Song, Mao Zheng, Xuan Luo Tencent Hunyuan nickmysong@tencent.com
@@ 1 Introduction
Recently, the needle-in-a-haystack benchmark [^2] https://github.com/gkamradt/LLMTest_NeedleInAHaystack has become a popular benchmark for evaluating whether LLMs have the capability of acquiring information from long documents. Specifically, this benchmark requires LLMs to precisely retrieve a specific sentence inserted at an arbitrary position within a long context, thereby assessing their capability to search for the sentence in long contexts. However, many newly released long-context LLMs have adopted the needle-in-a-haystack benchmark to evaluate their long-context processing capabilities and have achieved nearly perfect performance. This renders the needle-in-a-haystack benchmark insufficient to distinguish the differences between these models. Not only does this demonstrate the advancements in recent long-context LLMs, but it also indicates that the needle-in-a-haystack benchmark is too simplistic to further test their capabilities. Generally, acquiring information should be the most fundamental capability of long-text LLMs, which is the prerequisite for completing complex tasks. However, existing long-context benchmarks rarely focus on the multi-evidence collection ability of LLMs. Even when they do, the amount of evidence to be collected is relatively small and not proportional to the amount of evidence that a long document should contain.
To mitigate the shortcomings of existing benchmarks, in this paper, we propose a multi-evidence, position-aware, and scalable benchmark for evaluating long-context LLMs, named Counting-Stars. As the name suggests, the Counting-Stars refers to asking LLMs to count the numbers of stars from multiple sentences describing the number of stars counted by the little penguin inserted in the long context and then summarize into a specified answer. Through the Counting-Stars, we expect to evaluate the long context capabilities of multi-evidence searching and multi-evidence reasoning of LLMs. More specifically, the former focuses on testing the capability of LLMs to retrieve evidence at different positions within the long context, which can more clearly reflect the quality of long-context modeling. When collecting evidence, reasoning is often required to ensure that the evidence gathered supports the correct answer to the question. Therefore, the latter evaluates the LLM’s ability to filter out noise or incorrect information when retrieving information and the model’s reasoning ability at different positions within the long context. Generally speaking, the latter is definitely more challenging than the former. In other words, the former can be treated as making LLMs distinguish between long contexts and inserted sentences (similar to the needle-in-a-haystack benchmark), while the latter involves distinguishing evidence within each inserted sentence.
@@
Multi-evidence Searching
@@
Table 1: Prompt templates for the two counting tasks in the English version of Counting-Stars.
@@ 2 Counting-Stars (★)
Multi-evidence. In long-context scenarios, answering a question may require gathering substantial evidence from different positions within the lengthy context. Therefore, it is necessary to verify the ability of LLMs to collect a large amount of evidence from a long context all at once. Moreover, to the best of our knowledge, Counting-Stars is the first long-context benchmark to significantly increase the number of pieces of evidence (i.e., increase to $32$, $64$, $128$, $256$, $512$, and even $1024$).
@@
In general, the capacity for a long-context LLM to do human textual instructions largely depends on its multi-evidence searching ability. Moreover, an indispensable ability of LLMs extends beyond mere essential evidence collection to encompass reasoning based on the collected evidence. Therefore, the Counting-Stars mainly evaluates the long-context capability of LLMs from two perspectives, i.e., long-context multi-evidence searching and long-context multi-evidence reasoning. Expressly, this test can be understood as asking LLMs to find and remember all the sentences in the long text that describe the little penguin counting stars and organize them into a list to return the final answer. All sentences are prior inserted into a long text at the same interval. In addition, the used long text can be any text data that is not related to the sentences describing the little penguin counting stars, such as The Story of the Stone [^3] The Story of the Stone, is an 18th-century Chinese novel authored by Cao Xueqin, considered to be one of the Four Great Classical Novels of Chinese literature. for the Chinese version of the Counting-Stars and Paul Graham Essays [^4] The English context data used in this paper is similar to the needle-in-a-haystack. for the English version of the Counting-Stars. Next, we introduce the Counting-Stars in detail.
@@
2.1 Long-Context Multi-evidence Searching
Multi-evidence searching refers to the capability of distinguishing and collecting critical information framed within intricate and long textual data, which bottlenecks the performance of LLMs in synthesizing contextualized knowledge to execute various tasks, from answering multi-document questions to executing complex human instructions. Furthermore, maintaining a comprehensive and accurate grasp of the input text becomes increasingly challenging as the context length increases. Therefore, in the Counting-Stars test, the first task is to examine the multi-evidence searching ability of long-context LLMs, as illustrated in Table 1 (named as Long-Context Multi-evidence Searching). In multi-evidence searching, all sentences describing the little penguin counting stars are designed as "The little penguin counted {number1}". Here, {number1} indicates the number of stars the little penguin counted. Concretely, we randomly generated all the numbers of stars as {number1, number2, …} because we found that LLMs easily slack off if a sequence of numbers is continuous or regular. In this task, we hope that LLMs collect all the numbers of stars the little penguin counted and summarize them in a list.
@@ 2.3 Scalable Test Setting
P@32
Multi-evidence Searching (ZH) 0.697 0.663 0.775 0.807 0.788 0.698 0.682 0.606
Multi-evidence Searching (EN) 0.718 0.662 0.833 0.705 - - 0.389 0.559
P@32
@@
Table 3: The overall performance on the Counting-Stars-(32)-(Multi-evidence Reasoning).
Models Multi-evidence Searching (ZH)
@@
Table 4: The P@32 performance on the Chinese version of the Counting-Stars-(32)-(Multi-evidence Searching).
Models Multi-evidence Searching (EN) Multi-evidence Reasoning (ZH) Multi-evidence Reasoning (EN)
@@
Table 5: The P@32 performance on the English version of the Counting-Stars-(32)-(Multi-evidence Searching) as well as the Chinese and English versions of the Counting-Stars-(32)-(Multi-evidence Reasoning).
@@ 3.1 Baselines and Experimental Settings
Generally, evaluating text-based results is usually more complex, so the evidence to be collected in the Counting-Stars is all numerical, making it more straightforward to evaluate. For a piece of test data, the prediction results are evaluated starting from the first number of stars, that is, {number1}, {number2}, …, {numberM}. In this paper, we adopt P@N as the evaluation metric, which includes two modes: one where N=M, representing the counting times (here, M denotes the counting times); the other, referencing prior studies Yuan et al. (2020); Song et al. (2023a, b), adopts the P@M as the evaluation metric, where M denotes the total number of the retrieved results.
Specifically, for the Multi-evidence Searching task, if the results contain {number1}, it gets a score of $1$; if it doesn’t, it gets 00. Meanwhile, we construct a rule-based evaluation approach for the Multi-evidence Reasoning task, named P@32. For P@32, when the retrieved results contain only {number1}, the score is $1$; if it also contains {wrong number1}, the score is $0.5$; if the value only contains {wrong number1}, the score is $0.25$, and if both If not found, the score is 00.
Figure 2: Visualization of the results on the Chinese version of the Counting-Stars-32-(Multi-evidence Searching).
@@ 3.2 Overall Performance
Table 7 present the testing performance of GPT-4 Turbo, Claude3 Opus, Gemini 1.5 Pro, GLM-4, Moonshot-v1 on the Chinese and English versions of the Counting-Stars-(32)-(Multi-evidence Searching) and Counting-Stars-(32)-(Multi-evidence Reasoning) tests. Overall, Claude3 Opus achieves the best performance on the Chinese version of the Counting-Stars-(32)-(Multi-evidence Searching), Gemini 1.5 Pro obtains the best performance on the English version of the Counting-Stars-(32)-(Multi-evidence Searching) and the Chinese version of the Counting-Stars-(32)-(Multi-evidence Reasoning), and GPT-4 Turbo obtains the best performance on the English version of the Counting-Stars-(32)-(Multi-evidence Reasoning). Although these long-context LLMs have achieved nearly perfect performance on the needle-in-a-haystack task, they still perform poorly on the Counting-Stars test, which indicates that the needle-in-a-haystack is too simple to truly show the capabilities of LLMs in processing long texts.
The multi-evidence reasoning task necessitates that LLMs engage in acquiring and reasoning multiple pieces of evidence simultaneously, which is more complex than the multi-evidence searching task. This task requires LLMs to sift through and exclude inaccurate evidence while gathering information from a long context to answer questions. As indicated by the data in Table 7, each LLM performs not well enough. Notably, in contrast to GPT-4 Turbo and Claude3 Opus, Gemini 1.5 Pro stands out for having gathered virtually no incorrect information, as shown in Figure 3.
From Table 4 and Table 5, it can be observed that all LLMs are capable of achieving higher scores in short-context scenarios, which confirms that the Counting-Stars is reasonable and can be accomplished by LLMs. However, as the context length increases, the performance of all models shows a downward trend. Among them, GPT-4 Turbo’s performance is relatively stable. In addition, GLM-4 has obtained surprising results under the 32K context length of the Chinese version of the Counting-Stars-(32)-(Multi-evidence Searching).
@@ 4.1 The Length-Stability Dilemma
Based on the experiments in this paper, we are not yet able to determine the specific reasons behind this phenomenon; identifying these reasons is a goal that the next version of Counting-Stars aims to achieve. We consider the most intuitive explanation to be that the long-context capabilities of LLMs are still relatively weak, so when resources are limited, some stability must be sacrificed. Addressing this issue could help researchers better analyze and enhance the long-context modeling capabilities of LLMs, benefiting specific NLP tasks such as multi-document question answering. Moreover, stability refers to the understanding and reasoning abilities of LLMs when handling different long contexts, which is more crucial than merely the length of context processing.
@@
Models GPT-4 Turbo 1106 Gemini 1.5 Pro Claude3 Opus
F1@32 F1@M F1@32 F1@M F1@32 F1@M
Multi-evidence Searching (ZH) 0.808 0.808 0.866 0.866 0.887 0.889
Multi-evidence Searching (EN) 0.789 0.790 0.904 0.905 0.784 0.784
Multi-evidence Reasoning (ZH) 0.597 0.601 0.719 0.719 0.606 0.634
Multi-evidence Reasoning (EN) 0.757 0.769 0.373 0.404 0.457 0.468
Table 7: The F1@32 and F1@M performance on the Counting-Stars.
@@ 5 Related Work
The Counting-Stars evaluates the capabilities of multi-evidence searching and reasoning of LLMs, which should be more noteworthy in the long context modeling of LLMs, as reflected in tasks such as multi-document question answering and summarization. Concretely, the former primarily evaluates the capability of LLMs to collect multiple pieces of evidence simultaneously (distinguishing between long context and inserted sentences), while the latter tests the ability of LLMs to gather and reason various pieces of evidence at the same time correctly, that is, reasoning is required when collecting information (distinguishing between correct and incorrect evidence in inserted sentences). To the best of our knowledge, the Counting-Stars is the first scalable long-context benchmark to ask LLMs to simultaneously differentiate between correct and incorrect evidence in each inserted sentence.
@@ 6 Conclusion
We introduce a multi-evidence, position-aware, and scalable benchmark for evaluating the capabilities of the multi-evidence collection of long-context LLMs via two counting-based tasks. Utilizing the Counting-Stars benchmark, we conduct intriguing analyses on the behavior of LLMs, including the length-stability dilemma and the absence of the "lost-in-the-middle" phenomenon. Our analysis provides valuable insights into how LLMs handle long contexts, which can inform and guide future research endeavors.
7 Acknowledgments
We thank the three anonymous reviewers for carefully reading our paper and their insightful comments and suggestions.
@@ 8 Limitations
In the future, we will evaluate more famous LLMs on the Counting-Stars, such as Llama3888https://github.com/meta-llama/llama3, Mistral [^9] https://mistral.ai/news/la-plateforme/, Llama2 Touvron et al. (2023), Mixtral Jiang et al. (2024), Command-R [^10] https://docs.cohere.com/docs/command-r#model-details, LongLoRA Chen et al. (2024), LongAlpacaChen et al. (2024), LWM Liu et al. (2024), QWen Bai et al. (2023a), DeepSeek-V2 DeepSeek-AI (2024), etc.
@@
Initially, the Counting-Stars is designed to require LLMs to count the total number of stars in all sentences inserted in the long context, which aims to test the multi-evidence searching of LLMs from a long dependency perspective. However, we find that if LLMs are required to calculate the total number of stars, they usually perform poorly. Specifically, we analyze the reasons for the bad performance, which mainly include three points: